Please solve this question,
1m3 of olive oil has a mass of 920kg.
What is the volume of 460g of oil?
Regards

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

460 gm of olive oil is 0.46 kg
0.46 kg is 0.0005 of 920 kg.
1 m^3=1,000,000 cm^3 (100*100*100)
5*10^(-4)(10^6)=5*10^2 cm^3 or 500 cm^3. ANSWER.
This makes sense, because the density is 0.92 gm/cm^3, which is slightly less than water and consistent with an oil.

Answer by Alan3354(69443)

Half the mass --> half the volume.
================
Jeez.
-----------
Questions are not solved. Questions are answered.